1.     In a survey of 1500 households, it is found that 47% of them have a high-definition television (based on data from the consumer electronics association). Use a 0.01 significance level to test the claim that fewer than half of all households have a high-definition television.  Is the result from a few years ago likely to be valid today?



1
Expert's answer
2021-11-18T13:27:30-0500

"H_0: p_0 = 0.5 \\

H_1: p_0 < 0.5 \\

α=0.01 \\

N= 1500 \\

p = 0.47"

Test-statistic

"Z = \frac{p-p_0}{\sqrt{\frac{p_0(1-p_0)}{N}}} \\

Z = \frac{0.47-0.5}{\sqrt{\frac{0.5(1-0.5)}{1500}}} \\

Z = \frac{-0.03}{\sqrt{\frac{0.5 \times 0.5}{1500}}} \\

Z = -2.324"

At α=0.01 the critical value for a left-tailed test is "Z_c = -2.326"

Since "Z>Z_c" the null hypothesis is accepted.

There is NOT enough evidence to claim that fewer than half of all households have a high-definition television at 0.01 level of significance.

If the data and claim are the same in few years ago, the results likely to be valid today.